OpenStudy (anonymous):

how do you solve x^2-2xy=6 by implicit differentiation?

OpenStudy (anonymous):

\[= 2x -2yy'\]\[y' = -2x/2y \] \[y' = -\]

OpenStudy (anonymous):

y' = -x/y *

OpenStudy (anonymous):

correct me if I'm wrong ^_^

OpenStudy (anonymous):

dy/dx(x^2-2xy=6) = 2x(dx)-(2y(dx)+2x(dy)=0 =2x(dx)-2y(dx)-2x(dy)=0 =(dx)(2x-2y)=2x(dy) =(2x-2y)=2x(dy/dx) (dy/dx)=(2x-2y)/(2x) (dy/dx)=(x-y)/(x)

OpenStudy (anonymous):

lol, right! I forgot abt the 2xy, arman's answer is correct ^_^

OpenStudy (anonymous):

the book says the answer is 1- y/x

OpenStudy (anonymous):

yeah same thing

OpenStudy (anonymous):

hmm, let me write it down on paper

OpenStudy (anonymous):

(x-y)/(x) =(x/x)-(y/x) =1-(y/x)

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

did you understand it lefty? :)

OpenStudy (anonymous):

lots of use of distribution, and dont forget your product rule! :O

OpenStudy (anonymous):

you got that right, how can I forget such thing >_<

OpenStudy (anonymous):

yup, your above walkthrough i will figure out. thank you guys

OpenStudy (anonymous):

thanks to arman :) good luck!

OpenStudy (anonymous):

aha no worries, i messed up my quotient rule for the first 50+ pages of problem sets i did, i was devastated when i found out

OpenStudy (anonymous):

yep yep good lucck

OpenStudy (anonymous):

yeah it makes it even tougher my professor is a Romanian immigrant